# Invariants of Superalgebras as Complex Integrals

### v1

@@ Abstract
In [1] we defined integrals that approximated the Poincaré series of the invariants and cocomiants of the general linear Lie supergroup or superalgebra. Budzik suggested in [3] a way to adapt this method to get the exact Poincaré series. The purpose of this paper is to prove that Budzik’s ideas are correct. As a consequence we prove that the Poincaré series are rational functions.
@@ 1 Budzik’s Conjecture
The basic objects of that form the subject of this paper can be described in several different ways. The proof of our main theorem will mosty focus on their definition based on the character theory of the symmetic group, but in order to provide motivation and context, we also mention some of the connections to trace identities and to invariant theory.
@@
Sometimes we will write $m_{\lambda}(k,\ell)$ if we need to specify the $k$ and $\ell$. We associate to $m_{\lambda}$ the Poincaré series $P(k,\ell;n)=\sum m_{\lambda}s_{\lambda}(t_{1},\ldots,t_{n})$, where $s_{\lambda}$ is the Schur function, and more generally,
@@
for all $a_{1},\ldots,a_{n}\in M_{k,\ell}$ and all $g\in U(k|\ell)$. The algebra of invariant functions on $M_{k,\ell}^{n},$ polynomial in the entries, also has an $n$-fold grading and also has Poincaré series $P(k,\ell,n)$. In fact it is generated by the trace polynomials.
The series $P(k,\ell;n,m)$ also has a natural interpretation. We refer the interested reader to [1]. We also mention that we learned from [3] that these series have signifigance in physics: “These arise in physics e.g. as free partition functions of $U(N|M)$ gauge theories or supersymmetric indices.”
@@
Likewise, setting $P^{\prime}(k,\ell;n)$ equal to $\sum m_{\lambda}^{\prime}S_{\lambda}(t_{1},\ldots,t_{n})$, Budzik’s conjecture would imply that $P^{\prime}(k,\ell;n)=P(k,\ell;n)-P(k-1,\ell-1;n)$ and so
@@
More generally, setting $P^{\prime}(k,\ell;n,m)$ equal to $\sum m_{\lambda}^{\prime}HS(t_{1},\ldots,t_{n};u_{1},\ldots,u_{m})$ we would have $P^{\prime}(k,\ell;n,m)=P(k,\ell;n,m)-P(k-1,\ell-1k;n)$ and
@@
$\oint_{|y_{i}|=1}\oint_{|x_{i}|=1+\epsilon}\frac{\prod_{i}\prod_{z\in Z_{1}}(1+zt_{i})}{\prod_{i}\prod_{z\in Z_{0}}(1-zt_{i})\prod_{i}}\Delta\frac{dx_{1}}{x_{1}}\wedge\cdots\wedge\frac{dy_{\ell}}{y_{\ell}}$
Before leaving this section we also mention that Budzik computed $P^{\prime}(k,\ell;1)$ for all $k$ and $\ell$, namely,
$P^{\prime}(k,\ell;1)=\frac{t^{k\ell}}{\prod_{i=1}^{k}(1-t^{i})\prod_{j=1}^{\ell}(1-t^{j})}.$
@@ 2 Properties of Hook Schur Functions
We fix two sets of variables $X=\{x_{1},\ldots,x_{k}\}$ and $Y=\{y_{1},\ldots,y_{k}\}$. For each partition $\lambda$ we let $HS_{\lambda}(X;Y)$ be the hook Schur function on $\lambda$. This is commonly defined as $\sum_{\mu\subseteq\lambda}s_{\lambda}(X)s_{\lambda/\mu}(Y)$, where $s$ is the ordinary Schur function. Some authors study the super Schur funcitons $s_{\lambda}(X/Y)$ instead of the hook Schur functions $HS_{\lambda}(X;Y)$. Each has its advantages, but they are essentially equivalent, since $HS_{\lambda}(X;-Y)=s_{\lambda}(X/Y)$.
@@
The second theorem is due to Józefiak and Pragacz, [4]. It can also be found in [5], 1.3 Ex. 24. To each partition $\lambda$ we say $(i,j)\in\lambda$ if $\lambda_{i}\leq j$, or, more geometrically, if the Young tableau associated to $\lambda$ has a box in row $i$, column $j$. Then $f_{\lambda}$ is defined to be the product $\prod(x_{i}+y_{j})$ over all $(i,j)\in\lambda$, with the convention that $x_{i}=0$ and $y_{j}=0$ for $i>k$ or $j>\ell$. For example, if $(k,\ell)=(2,1)$ and $\lambda=(2,1)$, then $f_{\lambda}=(x_{1}+y_{1})x_{1}^{2}(x_{2}+y)$. Józefiak and Pragacz’s theorem gives an alternate description of hook Schur functions.
@@
Let $\lambda$ be a typical partition, and define the partitions $\alpha$ and $\beta$, via $\alpha=(\lambda_{1}-\ell,\ldots,\lambda_{k}-\ell)$ and $\beta=(\lambda_{1}^{\prime}-k,\ldots,\lambda_{\ell}^{\prime}-k)$. Then $HS_{\lambda}(X;Y)=\prod_{i,j}(x_{i}+y_{j})s_{\alpha}(X)s_{\beta}(Y)$.
@@ 3 The Inner Product
If $\mu$ and $\nu$ are not both in $H^{\prime}(k,\ell)$, then $\langle HS_{\mu},HS_{\nu}\rangle=0$.
@@
Let $f=f(z_{1},\ldots,z_{n})$ and $g=g(z_{1},\ldots,z_{n})$ be Laurant polynomials with leading terms of degrees $a$ and $b$, respectively, with a<b1a<b-1. Let $|z|=r$ be a circle containing all of the roots of $g(z)$. Then $\oint_{|z|=r}f(z)/g(z)dz=0$.
@@
Since $\mu$ is not typical, $\mu_{k}\leq\ell-1$, so $f_{\mu}(X;Y)$ is of degree at most $\ell-1$ in $x_{k}$; also $\prod_{i<j}(1-x_{j}x_{i}^{-1})^{-1}$ will be of degree k1k-1 in $x_{k}$, so (8) will be of degree at most $\ell-k$ in $x_{k}$, and (6) will be of degree at most 2. Under a permutation $\sigma$, $f_{\lambda}(X;Y)/\prod(1-x_{j}x_{i}^{-1})$ will be of degree at most $\ell-k$ in $x_{\sigma(k)}$, and the product with (7) at most 2-2. ∎
@@
such that in $x^{\alpha}=x^{\alpha_{1}}\cdots x^{\alpha_{k}},$ at least one $\alpha_{i}\leq-2$. The integral with respect to the $x_{j}$ can be broken up into the product of integrals
$\oint\frac{x_{i}^{\alpha_{j}}}{\prod_{m}(x_{j}+y_{m})^{2}}\frac{dx_{j}}{x_{j}}.$
@@
Now that we have this inner product, we can reformulate the equations for $m_{\lambda}^{\prime}$, $P(k,\ell;n)$ and $P^{\prime}(k,\ell;n,m)$ in terms of it.
@@ 5 Concomitants
In [1] we studied not only the invariants of $U(k|\ell)$, but also the concomitants. Each of the three definitions of $m_{\lambda}$ have analogues to multiplicities we call $\bar{m}_{\lambda}$. In terms of tensor products $\sum\bar{m}_{\lambda}$ is the multiplicity of $\chi^{\lambda}$ in $\sum_{\mu\in H(k,\ell)}(\chi^{\mu}\otimes\chi^{\mu})\downarrow,$ where the arrow denotes inducing down from each $S_{n}$ to $S_{n-1}$. For the second one, we use $\bar{F}[X_{1},\ldots,X_{n}]$, defined to be the tensor product $F[X_{1},\ldots,X_{n}]\otimes TR(k,\ell;n)$. This algebra is $n$-graded and has Poincaré series $\sum\bar{m}_{\lambda}s_{\lambda}(t_{1},\ldots,t_{n})$, denoted $\bar{P}(k,\ell;n)$. And, for the third one, we use the algebra of $U(k|\ell)$-invariant functions $M_{k,\ell}^{n}\rightarrow M_{k,\ell}$. This turns out to be isomorphic to the previous, and so also has Poincaré series $\bar{P}(k,\ell;n)$.
@@
$\quad\prod_{i}\prod_{z\in Z_{1}}(1+zt_{i})^{-1}\prod_{i}\prod_{z\in Z_{0}}(1+zu_{i})^{-1}\sum z\rangle$
@@
For each $i$, $\frac{\partial}{\partial t_{i+1}}P^{\prime}(k,\ell,i+1)|_{t_{i+1}=0}=\bar{P}^{\prime}(k,\ell;i)$.
@@
$\sum\bar{m}_{\lambda}^{\prime}(k,\ell)s_{\lambda}=\sum\bar{m}_{\lambda}^{\prime}(k,\ell)s_{\lambda}(k,\ell)-\sum\bar{m}_{\lambda}^{\prime}(k-1,\ell-1)s_{\lambda}(k,\ell)$
and so each $\bar{m}_{\lambda}^{\prime}(k,\ell)=\sum\bar{m}_{\lambda}(k,\ell)-\sum\bar{m}(k-1,\ell-1)$, which in term implies $\bar{P}^{\prime}(k,\ell,n,m)=P(k,\ell;n,m)-P(k-1,\ell-1;n,m)$.
@@ 6 Algebraic Properties of the Poincaré Series
By Theorem 1, $P^{\prime}(k,\ell;n)$ is the integral of
@@
In (9) the degree in each $x_{i}$ is $n(k+\ell)-1$, so by Lemma 8, the integral of (9) times $(-x^{-1}y_{j})^{m}$ will be zero if $m\geq n(k+\ell)+1$. Letting $N$ equal to $n(k+\ell)+1$, we can replace each $\sum_{m=0}^{\infty}(m+1)(-x_{i}^{-1}y_{j})^{m}$ by $\sum_{m=0}^{N}(m+1)(-x_{i}^{-1}y_{j})^{m}$ without affecting the integral. Hence, $P^{\prime}(k,\ell;n)$ is a linear combination of (many) integrals with integrands of the form
@@
That such an integral is a nice rational function follows directly from Theorem 3.4 of [9], which in turn is based on Stanley’s Corollary 3.8 in [8]. Van Den Bergh’s theory associates directed graphs to integrals of this form. In our case, the graph would have vertices $v[x_{i}]$ and $v[y_{j}]$, corresponding to the variables $x_{1},\ldots,y_{\ell}$, and with an edge $e[x_{i},x_{j},t_{\alpha}]$ connecting $v[x_{i}]$ with $v[x_{j}]$ for each term $(1-x_{i}x_{j}^{-1}t_{\alpha})$ in the denominator, and an edge $e[y_{i},y_{j},t_{\alpha}]$ from $v[y_{i}]$ to $v[y_{j}]$ for each $(1-y_{i}y_{j}^{-1}t_{\alpha})$. Then his Theorem 3.4 not only states that the integral is a nice rational function, but identitifies the denominator has a product of terms of the form $1-t^{C}$, where $C$ is an oriented cycle in the graph, and where $t^{C}$ means the product of the $t_{\alpha}$ corresponding to edges in $C$. Since a cycle cannot have repeated vertices, and since edges connect only the $v[x_{i}]$ or the $v[y_{j}]$, but not both, $|C|\leq\max\{k,\ell\}$, and so the degree of $t^{C}$ is at most $\max\{k,\ell\}$.
@@
In terms of the graph theory, we would add to the previous graph edges $e[x_{i},y_{j},u_{\alpha}]$ for each $(1-x_{i}y_{j}^{-1}u_{\alpha})$ and $e[y_{i},x_{j},u_{\alpha}]$ for each $(1-x_{i}^{-1}y_{j}u_{\alpha})$. The effect is that cycles could now have length as great as $k+\ell$, the number of vertices, and the degree in the edges corresponding to the $u_{\alpha}$ must be even.

### v2

@@ Abstract
In [2] we defined integrals that approximated the Poincaré series of the invariants and concomiants of the general linear Lie supergroup or superalgebra. Budzik suggested in [4] a way to adapt this method to get the exact Poincaré series. The purpose of this paper is to prove that Budzik’s ideas are correct. As a consequence we prove that the Poincaré series are rational functions.
@@ 1 Budzik’s Conjecture
The basic objects that form the subject of this paper can be described in several different ways. The proof of our main theorem will mosty focus on their definition based on the character theory of the symmetic group, but in order to provide motivation and context, we also mention some of the connections to trace identities and to invariant theory.
@@
Sometimes we will write $m_{\lambda}(k,\ell)$ if we need to specify the $k$ and $\ell$. We associate to $m_{\lambda}$ the Poincaré series
$P(k,\ell;n)=\sum m_{\lambda}s_{\lambda}(t_{1},\ldots,t_{n}),$
where $s_{\lambda}$ is the Schur function, and more generally,
@@
for all $a_{1},\ldots,a_{n}\in M_{k,\ell}$ and all $g\in U(k|\ell)$. The algebra of invariant functions on $M_{k,\ell}^{n},$ polynomial in the entries, also has an $n$-fold grading and also has Poincaré series $P(k,\ell,n)$, the same as above. In fact it is generated by the trace polynomials.
The series $P(k,\ell;n,m)$ also has a natural interpretation, see [1]. The algebra of $(k+\ell)\times(k+\ell)$ matrices with entries in $E$ has a $\mathbb{Z}_{2}$-grading in which $M_{k,\ell}$ is the degree 0 part. The degree 1 part would consist of matrices $\left(\begin{smallmatrix}A&B\\ C&D\end{smallmatrix}\right)$,with dimensions as above, but in which $A$ and $D$ have degree 1 entries and $B$ and $C$ have degree 0 ones. This algebra has a supertrace function, defined on a homogeneous matrix as above as $tr(A)\mp tr(D)$, depending on whether the matrix is degree 0 or 1. Letting $F[X_{1},\ldots,X_{n},Y_{1},\ldots,Y_{m}]$ be the algebra generated by $n$ generic degree 0 matrices and $m$ generic degree 1 matrices, we can then let $STR(k,\ell;n,m)$ be the algebra generated by the supertraces. This algebra has an $(n+m)$-fold grading and its Poincaré series is $P(k,\ell;n,m)$.
We also mention that we learned from [4] that these series have signifigance in physics: “These arise in physics e.g. as free partition functions of $U(N|M)$ gauge theories or supersymmetric indices.”
@@
Likewise, since $P^{\prime}(k,\ell;n)$ is equal to $\sum m_{\lambda}^{\prime}S_{\lambda}(t_{1},\ldots,t_{n})$, Budzik’s conjecture would imply that $P^{\prime}(k,\ell;n)=P(k,\ell;n)-P(k-1,\ell-1;n)$ and so
@@
More generally, since $P^{\prime}(k,\ell;n,m)$ is equal to $\sum m_{\lambda}^{\prime}HS_{\lambda}(t_{1},\ldots,t_{n};u_{1},\ldots,u_{m})$ we would have $P^{\prime}(k,\ell;n,m)=P(k,\ell;n,m)-P(k-1,\ell-1;n,m)$ and
@@
$\oint_{|y_{i}|=1}\oint_{|x_{i}|=1+\epsilon}\frac{\prod_{i}\prod_{z\in Z_{1}}(1+zt_{i})}{\prod_{i}\prod_{z\in Z_{0}}(1-zt_{i})}\Delta\frac{dx_{1}}{x_{1}}\wedge\cdots\wedge\frac{dy_{\ell}}{y_{\ell}}$
@@ 2 Properties of Hook Schur Functions
We fix two sets of variables $X=\{x_{1},\ldots,x_{k}\}$ and $Y=\{y_{1},\ldots,y_{k}\}$. For each partition $\lambda$ we let $HS_{\lambda}(X;Y)$ be the hook Schur function on $\lambda$. This is commonly defined as $\sum_{\mu\subseteq\lambda}s_{\mu}(X)s_{\lambda/\mu}(Y)$, where $s$ is the ordinary Schur function. Some authors study the super Schur funcitons $s_{\lambda}(X/Y)$ instead of the hook Schur functions $HS_{\lambda}(X;Y)$. Each has its advantages, but they are essentially equivalent, since $HS_{\lambda}(X;-Y)=s_{\lambda}(X/Y)$.
@@
The second theorem is due to Józefiak and Pragacz, [5]. It can also be found in [6], 1.3 Ex. 24. To each partition $\lambda$ we say $(i,j)\in\lambda$ if $\lambda_{i}\leq j$, or, more geometrically, if the Young tableau associated to $\lambda$ has a box in row $i$, column $j$. Then $f_{\lambda}$ is defined to be the product $\prod(x_{i}+y_{j})$ over all $(i,j)\in\lambda$, with the convention that $x_{i}=0$ and $y_{j}=0$ for $i>k$ or $j>\ell$. For example, if $(k,\ell)=(2,1)$ and $\lambda=(2,1)$, then $f_{\lambda}=(x_{1}+y_{1})x_{1}(x_{2}+y)$. Józefiak and Pragacz’s theorem gives an alternate description of hook Schur functions.
@@
Let $\lambda$ be a typical partition, and define the partitions $\mu$ and $\nu$, via $\mu=(\lambda_{1}-\ell,\ldots,\lambda_{k}-\ell)$ and $\nu=(\lambda_{1}^{\prime}-k,\ldots,\lambda_{\ell}^{\prime}-k)$, see Figure 1. Then $HS_{\lambda}(X;Y)=\prod_{i,j}(x_{i}+y_{j})s_{\mu}(X)s_{\nu}(Y)$.
@@ 3 The Inner Product
If $\mu$ and $\nu$ are not both typical, then $\langle HS_{\mu},HS_{\nu}\rangle=0$.
@@
Let $f=f(z_{1},\ldots,z_{n})$ and $g=g(z_{1},\ldots,z_{n})$ be Laurant polynomials with leading terms of degrees $a$ and $b$, respectively, with a<b1a<b-1. Let $|z|=r$ be a circle containing all of the roots of $g(z)$ in its interior. Then $\oint_{|z|=r}f(z)/g(z)dz=0$.
@@
Since $\mu$ is not typical, $\mu_{k}\leq\ell-1$, so $f_{\mu}(X;Y)$ is of degree at most $\ell-1$ in $x_{k}$; also $\prod_{i<j}(1-x_{j}x_{i}^{-1})^{-1}$ will be of degree k1k-1 in $x_{k}$, so (8) will be of degree at most $\ell-k$ in $x_{k}$. Under a permutation $\sigma$, $f_{\lambda}(X;Y)/\prod(1-x_{j}x_{i}^{-1})$ will be of degree at most $\ell-k$ in $x_{\sigma(k)}$, and the product with (7) at most 2-2. ∎
@@
such that for at least one $i$, $x_{i}^{\alpha_{i}}\prod_{j}(x_{i}+y_{j})^{-2}$the degree in $x_{i}$ is less than or equal to 2-2. The integral with respect to the $x_{j}$ can be broken up into the product of integrals
$\oint\frac{x_{j}^{\alpha_{j}}}{\prod_{m}(x_{j}+y_{m})^{2}}\frac{dx_{j}}{x_{j}}.$
@@ 4 Proof of Budzik’s Conjecture
Note that using Theorem 6 we can directly compute the formula for $P^{\prime}(k,\ell;n,m)$:
$P^{\prime}(k,\ell;n,m)$ $=\sum m_{\lambda}HS_{\lambda}(T,U)$
$=\sum\langle HS_{\lambda}(Z_{0};Z_{1}),1\rangle HS_{\lambda}(T;U)$
$=\langle\sum HS_{\lambda}(Z_{0};Z_{1})HS_{\lambda}(T;U),1\rangle$
The sum is the integrand of Theorem  by the Berele-Remmel Theorem, Theorem 6.
@@ 5 Concomitants
In [2] we studied not only the invariants of $U(k|\ell)$, but also the concomitants. Each of the three definitions of $m_{\lambda}$ have analogues to multiplicities we call $\bar{m}_{\lambda}$. In terms of tensor products $\sum\bar{m}_{\lambda}$ is the multiplicity of $\chi^{\lambda}$ in $\sum_{\mu\in H(k,\ell)}(\chi^{\mu}\otimes\chi^{\mu})\downarrow,$ where the arrow denotes inducing down from each $S_{n}$ to $S_{n-1}$. For the second one, we use $\bar{F}[X_{1},\ldots,X_{n}]$, defined to be the tensor product $F[X_{1},\ldots,X_{n}]\otimes TR(k,\ell;n)$. This algebra is $n$-graded and has Poincaré series $\sum\bar{m}_{\lambda}s_{\lambda}(t_{1},\ldots,t_{n})$, denoted $\bar{P}(k,\ell;n)$. And, for the third one, we use the algebra of $U(k|\ell)$-invariant functions $M_{k,\ell}^{n}\rightarrow M_{k,\ell}$. This turns out to be isomorphic to the previous, and so also has Poincaré series $\bar{P}(k,\ell;n)$. Likewise for $\bar{P}(k,\ell;n,m)$.
@@
$\quad\prod_{i}\prod_{z\in Z_{1}}(1+zt_{i})^{-1}\prod_{i}\prod_{z\in Z_{0}}(1+zu_{i})^{-1}\sum z,1\rangle$
@@
For each $i$, $\frac{\partial}{\partial t_{i+1}}P^{\prime}(k,\ell;i+1)|_{t_{i+1}=0}=\bar{P}^{\prime}(k,\ell;i)$.
@@
$\sum\bar{m}_{\lambda}^{\prime}(k,\ell)s_{\lambda}=\sum\bar{m}_{\lambda}^{\prime}(k,\ell)s_{\lambda}-\sum\bar{m}_{\lambda}^{\prime}(k-1,\ell-1)s_{\lambda}.$
Since the Schur functions are linearly independent, this implies that $\bar{m}_{\lambda}^{\prime}(k,\ell)=\sum\bar{m}_{\lambda}(k,\ell)-\sum\bar{m}(k-1,\ell-1)$, which in term implies $\bar{P}^{\prime}(k,\ell,n,m)=P(k,\ell;n,m)-P(k-1,\ell-1;n,m)$.
@@ 6 Algebraic Properties of the Poincaré Series
By Theorem 4, $P^{\prime}(k,\ell;n)$ is the integral of the product of
@@
In (9) the degree in each $x_{i}$ is $n(k+\ell)-1$, so by Lemma 8, the integral of (9) times $(-x^{-1}y_{j})^{m}$ will be zero if $m\geq n(k+\ell)+1$. Letting $N$ equal to $n(k+\ell)+1$, we can replace each $\sum_{m=0}^{\infty}(m+1)(-x_{i}^{-1}y_{j})^{m}$ by $\sum_{m=0}^{N}(m+1)(-x_{i}^{-1}y_{j})^{m}$ without effecting the integral. Hence, $P^{\prime}(k,\ell;n)$ is a linear combination of (many) integrals with integrands of the form
@@
That such an integral is a nice rational function follows directly from Theorem 3.4 of [13], which in turn is based on Stanley’s Corollary 3.8 in [12]. Van Den Bergh’s theory associates directed graphs to integrals of this form. In our case, the graph would have vertices $v[x_{i}]$ and $v[y_{j}]$, corresponding to the variables $x_{1},\ldots,y_{\ell}$, and with an edge $e[x_{i},x_{j},t_{\alpha}]$ connecting $v[x_{i}]$ with $v[x_{j}]$ for each term $(1-x_{i}x_{j}^{-1}t_{\alpha})$ in the denominator, and an edge $e[y_{i},y_{j},t_{\alpha}]$ from $v[y_{i}]$ to $v[y_{j}]$ for each $(1-y_{i}y_{j}^{-1}t_{\alpha})$. Then his Theorem 3.4 not only states that the integral is a nice rational function, but identitifies the denominator has a product of terms of the form $1-t^{C}$, where $C$ is an oriented cycle in the graph, and where $t^{C}$ means the product of the $t_{\alpha}$ corresponding to edges in $C$. Since a cycle cannot have repeated vertices, and since edges connect only the $v[x_{i}]$ to each other or the $v[y_{j}]$ to each other, $|C|\leq\max\{k,\ell\}$, and so the degree of $t^{C}$ is at most $\max\{k,\ell\}$.
@@
In terms of the graph theory, we would add to the previous graph edges $e[x_{i},y_{j},u_{\alpha}]$ for each $(1-x_{i}y_{j}^{-1}u_{\alpha})$ and $e[y_{i},x_{j},u_{\alpha}]$ for each $(1-x_{i}^{-1}y_{j}u_{\alpha})$. The effect of these additional edges would be that cycles could now have length as great as $k+\ell$, the number of vertices, and the degree in the edges corresponding to the $u_{\alpha}$ must be even.
7 Computation of $P^{\prime}(k,\ell;0,1)$
In [4] Budik computed $P^{\prime}(k,\ell,1,0)$, the Poincaré series of the ring generated by the traces of a single generic element of $M_{k,\ell}$:
$P^{\prime}(k,\ell;1)=\frac{t^{k\ell}}{\prod_{i=1}^{k}(1-t^{i})\prod_{j=1}^{\ell}(1-t^{j})},$ (12)
Also computed by Molev and Mukhin in [7], Orellana and Zabrocki in [8]. A different formula for $P(k,\ell;n)$ was given by Sergeev and Veselov in [11]. We give yet another derivation based on Theorem 5.2 of [2], our proof being completely combinatorial.
Theorem 14 (=Theorem 5.2 of [2]).
$m_{(n)}(k,\ell)$ equals the number of partitions of $n$ in the hook $H(k,\ell)$; and $m_{(1^{n})}(k,\ell)$ equals the number of self-conjugate partitions of $n$ in $H(k,\ell)$.
It follows that $m^{\prime}_{(n)}(k,\ell)=m_{(n)}(k,\ell)-m_{(n)}(k-1,\ell-1)$ equals the number of partitions in $H(k,\ell)$, but not in $H(k-1,\ell-1)$. These are the typical partitions, $H^{\prime}(k,\ell)$. A Young diagram of a typical partition consists of three parts: The $k\times\ell$ rectangle, a partition of height at most $k$ to the right of it, and the conjugate of a partition of height at most $\ell$ below it. See figure 1. Since the generating function for partitions of height at most $k$ is $\prod_{i=1}^{k}(1-x^{i})^{-1}$, (12) follows.
$k\times\ell$$\mu$$\nu^{\prime}$
Figure 1: A Typical Partition
We now turn our attention to $P^{\prime}(k,\ell;0,1)$. We will take $k\geq\ell$. This is the Poincaré series of the supertrace ring generated by one matrix of the form $\left(\begin{smallmatrix}A&B\\ C&D\end{smallmatrix}\right)$, where $A$ is a $k\times k$ matrix, $D$ is an $\ell\times\ell$ matrix, $A$ and $D$ have generic odd Grassmann entries, and $B$ and $C$ have entries commuting indeterminants.
As above,
$m_{(1^{n})}^{\prime}(k,\ell)=m_{(1^{n})}(k,\ell)-m_{(1^{n})}(k-1,\ell-1),$
which, by Theorem 14, equals the number of self-conjugate partitions of $n$ in the hook $H(k,\ell)$, but not in the hook $H(k-1,\ell-1)$.
If a self-conjugate partition $\lambda$ lies in the hook $H(k,\ell)$, it will also lie in $H(\ell,k)$, meaning that $\lambda_{k+1}\leq\ell$ and $\lambda_{\ell+1}\leq k$. In terms of Young diagrams, such a $\lambda$ can be broken into three parts.
Definition.
Given a partition $\lambda$ we define $\lambda_{0}$ to be partition $\lambda\cap(k^{k})$, namely, the partition whose Young diagram is the intersection of the Young diagram of $\lambda$ with that of the square $(k^{k})$; $\mu=\mu(\lambda)$ to be the partition to the right of $\lambda_{0}$; and $\nu=\nu(\lambda)$ to be the conjugate of the partition below it, see Figure 2.
Note that $\lambda$ is self-conjugate if and only if $\lambda_{0}$ is self-conjugate and $\mu(\lambda)=\nu(\lambda)$.
Lemma 15.
The correspondence $\lambda\rightarrow(\lambda_{0},\mu(\lambda))$ is a one-to-one correspondence between self-conjugate partitions in $H^{\prime}(k,\ell)$ and pairs of partitions, where $\mu$ is of height at most $\ell$ and $\lambda_{0}$ is a self-conjugate partition containing $(k^{\ell},\ell^{k-\ell})$ and contained in $(k^{k})$.
Proof.
If $\lambda\notin H(k-1,\ell-1)$ then $\lambda_{k}\geq\ell$ and, since $\lambda$ is self-conjugate, $\lambda_{\ell}\geq k$. The latter implies that the first $\ell$ rows of $\lambda_{0}$ must equal $k$, and the former that the first $\ell$ columns also equal $k$. Hence, $\lambda_{0}$ contains $(k^{\ell},\ell^{k-\ell})$.
The condition $\lambda\in H(k,\ell)\cap H(\ell,k)$ implies that $\mu$ has height at most $\ell$, and the conditions $\lambda_{\ell}\geq k$ and $\lambda_{k}\geq\ell$ imply that any partition of height at most $\ell$ is possible. ∎
$\lambda_{0}$$\nu$$\mu\ell\ellkkkk$
Figure 2: Construction of $\lambda_{0}$, $\mu$ and $\nu$
Theorem 16.
$P^{\prime}(k,\ell;0,1)=u^{2k\ell-\ell^{2}}\prod_{i=1}^{k-\ell}(1+u^{2i-1})\prod_{i=1}^{\ell}(1-u^{2i})^{-1}$.
Proof.
The generating function for partitions of height at most $\ell$ is $\prod_{i=1}^{\ell}(1-u^{i})^{-1}$, so the generating function for equal pairs of them is $\prod_{i=1}^{\ell}(1-u^{2i})^{-1}$. As for $\lambda_{0}$, $\lambda_{0}$ is made up of two parts: $(k^{\ell},\ell^{k-\ell})$ and a self-conjugate partition contained in the lower right $(k-\ell)\times(k-\ell)$ square. The former has degree $k\ell+\ell(k-\ell)=2k\ell-\ell^{2}$, and the latter has generating function $\prod_{i=0}^{k-\ell}(1+u^{2i-1}).$
Budzik related the formula for $P^{\prime}(k,\ell;n)$ to various important combinatorial identities. The same came be done for $P^{\prime}(k,\ell;0,1)$. The simplest is based on $k=\ell$, in which case
$P^{\prime}(k,k;0,1)=\frac{u^{k^{2}}}{[u^{2}]_{k}},$
the generating function for self-conjugate partitions in $H^{\prime}(k,k)$. Using $P(k,k;0,1)=\sum_{n=0}^{k}P^{\prime}(n,n;0,1)$ and letting $k$ go to infinity we get the well-known identity
$\prod_{n=1}^{\infty}(1+u^{2n-1})=\sum_{k=0}^{\infty}\frac{u^{k^{2}}}{[u^{2}]_{k}}$
since $P(\infty,\infty;0,1)$ is the generating function for all self-conjugate partitions.
Similarly, using the identity $P(n+\ell,\ell;0,1)=\sum_{i=0}^{\ell}P^{\prime}(n+i,i;0,1)$ and letting $\ell$ go to infinity, we get
$\prod_{i=1}^{\infty}(1+u^{2i-1})=\sum_{i=0}^{\infty}\frac{u^{i(2n+i)}\prod_{t=1}^{i}(1+u^{2t-1})}{[u^{2}]_{i}}.$